# On-board Satellite Image Classification for Earth Observation: A Comparative Study of Pre-trained Vision-Transformer Models

### v1

@@ III-B1 Training from Scratch
  • CNN: A basic convolutional neural network designed for IC, consisting of two convolutional layers followed by batch normalization, ReLU activation, and max pooling, with fully connected layers at the end for classification.
  • ResNet: A smaller version of the ResNet framework, termed ResNet-14, with 14 layers, including two residual blocks in each of the three hidden layers, optimizing feature extraction and gradient flow through the network.
@@
  • Small ViT [27]: A Vision Transformer model designed for efficiency, using a small patch tokenization process and a lightweight transformer with self-attention mechanisms to classify images effectively.
@@ III-B2 Pretrained ViT Models
  • EfficientViT-M2 [28]: An efficient vision transformer combining convolutional layers with local window attention mechanisms optimized for balancing performance and computational efficiency with approximately 4 million parameters.
@@
  • EfficientNet-B2 [31]: A highly efficient convolutional neural network using depthwise separable convolutions and squeeze-and-excitation layers to minimize parameter usage while maximizing performance.
@@
  • SwinTransformer [34]: A hierarchical vision transformer that divides images into non-overlapping patches, utilizing shifted window-based self-attention to capture both local and global context.
  • Vision Transformer (ViT) [35]: A transformer model that divides images into patches, processing them through multiple transformer layers with self-attention to achieve high-capacity IC.
@@ III-C Evaluation Metrics
To comprehensively evaluate the performance of our multiclass classification model across 10 classes, we employ three key metrics: accuracy, precision, and recall (sensitivity) [36]. These metrics are calculated for each class individually and then aggregated using macro-averaging as follows,
[…]
$\text{Precision}=\frac{1}{N}\sum_{i=1}^{N}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FP}_{i}}$
$\text{Recall}=\frac{1}{N}\sum_{i=1}^{N}\frac{\text{TP}_{i}}{\text{TP}_{i}+\text{FN}_{i}}$
By using macro-averaging, we ensure that each class is given equal importance, thereby providing a balanced evaluation of the model’s classification capabilities across the entire dataset. These macro-averaged evaluation metrics will be used to select the best models in the final analysis.
To assess the model’s robustness, we adopt a benchmarking approach for evaluating neural network robustness to common corruptions and perturbations [37]. The robustness score measures the degradation in accuracy caused by a specific perturbation, $\phi$. For a given model, $m$, let $A_{\text{clean}}$ represent the model’s accuracy on the original (clean) test dataset, and $A_{\phi}$ denote the accuracy of the model on the test set that has been modified with perturbation $\phi$. These perturbations can stem from natural environmental conditions encountered during deployment or adversarial modifications of the original test samples. Following the methodology outlined in [38], we define the robustness score of model $f$ with respect to perturbation $\phi$ as follows:
$R^{\phi}_{f}=\frac{A_{\phi}}{A_{\text{clean}}}$ (1)
A robustness score close to one indicates that the model is highly resistant to perturbation, demonstrating its ability to maintain performance in challenging conditions.
@@ IV Experimental Setup
The dataset was split into training and testing sets with a 70/30 ratio to ensure a balanced and comprehensive evaluation of the model’s performance.
For model optimization, different strategies were employed based on the model architecture. The CNN and ResNet models were optimized using the Stochastic Gradient Descent (SGD) optimizer [39], configured with a learning rate of 1e-3, momentum of 0.9, and a weight decay of 5e-4. In contrast, the ViT-based models, including the pre-trained variants, were optimized using the AdamW optimizer [40]. This setup also utilized a learning rate 1e-3 and a weight decay 5e-4, ensuring consistency across different model types. A step learning rate scheduler was applied to further enhance training efficiency with a step size of 7 epochs and a decay factor ($\gamma$) of 0.1.
The Compact Convolutional Transformer (CCT) [26] and SmallViT models [27] were trained using code adapted from the ViT-PyTorch repository, which provides reliable and optimized implementations for Vision Transformers. Other pre-trained models were sourced from Huggingface’s model hub or the official repositories in the respective papers. This approach ensured that the models were implemented consistently, with standardized architectures and preprocessing steps, allowing for a fair and direct comparison of their performance across the different experiments. 111 Available codes and pre-trained models:
@@
All experiments were conducted on a workstation equipped with an Intel(R) Xeon(R) W-11855M CPU, operating at 3.20GHz with 12 cores and supported by 64 GB of RAM. The models were trained and tested using an Nvidia RTX A5000 GPU with 48 GB of GDDR6 memory, providing sufficient computational resources for most tasks. This setup offered a robust and efficient environment for handling most models’ training and evaluation processes. However, due to their increased complexity and size, more extensive computational resources were necessary for the two larger pre-trained Vision Transformer models (ViT Large and ViT Huge). These models were parallelly trained on a high-performance computing workstation with the following specifications: an Intel Xeon w9-3475X CPU featuring 36 cores and 72 threads, a maximum turbo frequency of 4.8 GHz, and an 82.5 MB cache. The system was also equipped with 512 GB of DDR5 RAM, clocked at 4800 MHz, and dual NVIDIA RTX™ 6000 Ada Generation GPUs, each with 48 GB of GDDR6 memory. This advanced hardware configuration ensured efficient training of the larger models, leveraging parallel processing and extensive memory resources.
After identifying the best model with lower computational complexity and high performance, we further assess their robustness during inference onboard satellites. This analysis evaluates the models’ performance under noisy conditions, which is common in real-world satellite operations. Specifically, we introduce two types of noise into the inference test data: Gaussian noise and instrument noise represented by motion blur. For every kind of noise, we define five levels of severity, as illustrated in the image. This evaluation helps us understand how well the models can perform under varying noise levels, which is critical for reliable satellite-based IC, as shown in Fig. 2. The noise and severity level was implemented based on the code  [^2] https://github.com/hendrycks/robustness from [37].
Figure 2: Different noise levels.
@@
To further assess the robustness and generalization capabilities of the selected models, we applied a series of data augmentation techniques, as illustrated in Fig. 3. The augmentations include Standard Transformed, Rand Augment [41], Random Erasing [42], Rand(Augment+Erasing), and Strong Augmentation, which combine all four augmentations designed to simulate challenging variations in the input data. These transformations introduce different levels of distortion and occlusion, such as color shifts, random erasures, and severe augmentations, which mimic real-world variations that the models may encounter during satellite IC tasks. By evaluating the models on these augmented datasets, we aim to determine their ability to maintain performance when faced with altered or corrupted inputs, thereby ensuring their reliability in diverse and unpredictable conditions.
@@ V Results and Discussions
TABLE II: Summarization of model computation complexity, estimated total size, training, and inference times
Models Input Size Total Parameters Estimated Total Size (MB) FLOPs Training (s) Inference (s)
@@
Figure 4: Model complexity comparison.
The heatmap shown in Table II Fig. 4 provides a detailed comparison of the computational complexity of various models, with a particular emphasis on total parameters, estimated total size, FLOPs, training time, and inference time, all displayed on a logarithmic scale. Notably, the ViT-large and ViT-huge models stand out for their substantial complexity. They are characterized by high parameter counts, large memory footprints, and extensive computational operations (FLOPs), resulting in significantly longer training and inference times. In contrast, smaller models like EfficientViT-M2 and MobileViT2, despite being part of the Vision Transformer family, demonstrate much lower complexity across all metrics. These models are designed to be more lightweight, with fewer parameters and reduced FLOPs, making them more efficient in terms of both computational and memory resources. Compared to the pre-trained ViT models, EfficientViT-M2 and MobileViT2 balance performance and efficiency, providing a viable alternative for limited computational resources or faster processing times are crucial. This comparison underscores the diversity within Transformer-based models and highlights the importance of selecting models that align with the specific needs of a given application.
@@
TABLE III: Experiment results for the best model’s performance
@@
Based on the best run for each model in Table III, the summarized experimental results identify MobileViTV2, EfficientViT-M2, and SwinTransformer as the top-performing models. MobileViTV2 (Apple) achieves the highest overall performance, with a perfect training accuracy of 100%, leading to test accuracy, precision, and recall, each at 99.09%. SwinTransformer follows closely, boasting a training accuracy of 99.8% and a test accuracy, precision, and recall of 98.83%. EfficientViT-M2 ranks third, delivering impressive results with a training accuracy of 99.99% and test accuracy, precision, and recall, all at 98.76%. When comparing the top four large pre-trained models, the SwinTransformer notably outperforms its counterparts despite having a smaller parameter count of 86 million. It surpasses the ViT-base, ViT-large, and ViT-huge models, which do not achieve better results despite their larger parameter sizes. This highlights the effectiveness of the Swin Transformer’s design, which utilizes a hierarchical vision transformer architecture. Unlike conventional Vision Transformers (ViT), which rely on a global attention mechanism across the entire image, the SwinTransformer divides images into non-overlapping patches and hierarchically processes them. This method allows it to handle images at various scales and resolutions efficiently, resulting in better generalization and improved task performance. The superior performance of the SwinTransformer, even with fewer parameters, underscores the advantages of its architecture in effectively capturing both local and global image features, making it a more efficient and effective model compared to the larger, conventional attention-based ViT models. Explain why SwinTransformer is better than ViT
@@
From the results presented in Table II, and III, Fig. 4, and 5, we confirm that MobileViT and EfficientViT are the two best models for IC. These models excel due to their lower computational complexity, shorter training and inference times, and superior performance across all evaluation metrics. Notably, compared to state-of-the-art models, as shown in Table IV, the exceptional performance of EfficientViT-M2 and MobileViTV2 becomes evident. In the EuroSAT IC task, MobileViTV2 achieves the highest scores across all evaluation metrics, outperforming all other models in the comparison. EfficientViT-M2 also performs exceptionally well, securing the second-highest results. These findings highlight the effectiveness of both EfficientViT-M2 and MobileViTV2 in achieving top-tier performance with relatively lower computational complexity and faster training and inference times, as observed in previous experiments. This confirms that these two models are currently the best choices for high-performance IC tasks, providing a significant advantage over existing methods.
@@
Additionally, Fig. 6 shows the confusion matrices for EfficientViT-M2 and MobileViTV2, highlighting their exceptional performance in IC. Both models exhibit vital accuracy across all classes, with most predictions aligning closely with the true labels. In the confusion matrix for EfficientViT-M2, we observe that classes like AnnualCrop, Forest, PermanentCrop, and SeaLake are classified with nearly perfect accuracy, showing minimal misclassification. Similarly, the confusion matrix for MobileViTV2 shows even fewer errors, with several classes like Forest, PermanentCrop, and SeaLake being classified almost flawlessly. A few minor misclassifications are observed in both models, particularly in categories like Pasture and River, but these are relatively small compared to the overall high performance. These matrices underscore the effectiveness of both models in handling complex IC tasks, with MobileViTV2 slightly outperforming EfficientViT-M2 in terms of fewer misclassifications, further affirming its position as the top-performing model. Explain why EfficientViT and MobileViT are better than large ViT-based models
@@
TABLE IV: Performance comparison with the-state-of-the-art models
@@
The experiment results, as presented in Tables V and VI, reveal that applying data augmentation techniques during training and inference generally leads to a slight reduction in model performance when evaluated on clean test data. MobileViTV2 and EfficientViT-M2’s baseline performance—achieved without augmentation—yields the highest accuracy. However, among the various augmentation techniques tested, RandErasing for both models demonstrates the most effective results, providing performance metrics (98.73% for MobileViTV2 and 98.22% for EfficientViT-M2) that are closest to the baseline. This indicates that RandErasing introduces sufficient variability to enhance robustness without significantly compromising the models’ ability to perform well on clean data. Although there is a slight decrease compared to the baseline, using RandErasing offers a good balance, maintaining high accuracy while preparing the models to handle real-world scenarios with noisy inputs.
Figure 7: MobileViT (Top) and EfficientViT (Bottom) robustness with Gaussian noisy inference data
@@
Furthermore, the feasibility of deploying the EfficientViT-M2 model in practical applications, mainly onboard satellite systems, is underscored by its relatively low computational complexity and compact size. With an estimated total size of 38.19 MB and a computational complexity of 203.53 MFLOPs, EfficientViT-M2 offers a significant advantage over earlier models used in pioneering satellite missions. For instance, the $\Phi$-Sat-1 mission [11], which utilized the CloudScout model for image segmentation, required approximately 4.67 GFLOPs and had an estimated total size of 123.98 MB. In contrast, EfficientViT-M2 requires substantially less computational power and storage, making it a more efficient choice for similar tasks. Additionally, compared to the $\Phi$-Sat-2 mission [12], which employed a convolutional autoencoder with 81.2 MFLOPs for vision processing on an Intel Movidius Myriad 2 VPU, EfficientViT-M2’s slightly higher complexity is justified by its enhanced robustness and performance capabilities. While the $\Phi$-Sat-2 model took nearly 11 seconds to compress an image with a compression ratio of 8 on a VPU, EfficientViT-M2’s processing requirements are more aligned with modern GPU capabilities, completing complex tasks much faster and with greater accuracy. These factors make EfficientViT-M2 a powerful model for IC and a practical and feasible option for real-world deployment in resource-constrained environments like satellite missions.

### v2

@@ III-B1 Training from Scratch
  • CNN: A basic CNN designed for IC, consisting of two convolutional layers followed by batch normalization, ReLU activation, and max pooling, with fully connected layers at the end for classification.
  • ResNet-14: A smaller version of the ResNet framework, termed ResNet-14, with 14 layers, including two residual blocks in each of the three hidden layers, optimizing feature extraction and gradient flow through the network.
@@
  • Small ViT [35]: A small ViT model designed for efficiency, using a small patch tokenization process and a lightweight transformer with self-attention mechanisms to classify images effectively.
@@ III-B2 Pretrained ViT Models
  • EfficientViT-M2 [20]: An efficient ViT combining convolutional layers with local window attention mechanisms optimized for balancing performance and computational efficiency with approximately 4 million parameters.
@@
  • EfficientNet-B2 [37]: A highly efficient CNN using depthwise separable convolutions and squeeze-and-excitation layers to minimize parameter usage while maximizing performance.
@@
  • SwinTransformer [40]: A hierarchical ViT that divides images into non-overlapping patches, utilizing shifted window-based self-attention to capture both local and global context.
  • ViT [41]: A transformer model that divides images into patches, processing them through multiple transformer layers with self-attention to achieve high-capacity IC.
@@ III-C Evaluation Metrics
To comprehensively evaluate the performance of our multiclass classification model across 10 classes, we employ three key metrics: accuracy, precision, and recall (sensitivity) [42]. These metrics are calculated for each class individually and then aggregated using macro-averaging to assess the model’s performance as follows,
$\text{Accuracy}=\sum_{k=1}^{K}\frac{\text{TP}_{k}}{N},$
$\text{Precision}=\frac{1}{N}\sum_{i=1}^{K}N_{k}\frac{\text{TP}_{k}}{\text{TP}_{k}+\text{FP}_{k}},$
$\text{Recall}=\frac{1}{N}\sum_{i=1}^{K}N_{k}\frac{\text{TP}_{k}}{\text{TP}_{k}+\text{FN}_{k}},$
where $N$ is the total number of data points across all classes. $K$ is the total number of classes. $N_{k}$ is the number of data points in class $k$. $TP_{k}$ is True Positives, $FP_{k}$ is False Positives, $FN_{k}$ is False Negatives for class $k$, respectively. Using weighted precision and recall, we ensure that each class is given equal importance, thereby providing a balanced evaluation of the model’s classification capabilities across the entire dataset. These macro-averaged evaluation metrics will be used to select the best models in the final analysis.
To assess the model’s robustness, we adopt a benchmarking approach for evaluating NN robustness to common corruptions and perturbations [43]. The robustness score measures the degradation in accuracy caused by a specific perturbation, $\phi$. For a given model, $m$, let $A_{\text{clean}}$ represent the model’s accuracy on the original (clean) test dataset, and $A_{\phi}$ denote the accuracy of the model on the test set that has been modified with perturbation $\phi$. These perturbations can stem from natural environmental conditions encountered during deployment or adversarial modifications of the original test samples. Following the methodology outlined in [44], we define the robustness score of model $f$ with respect to perturbation $\phi$ as:
$R^{\phi}_{f}=\frac{A_{\phi}}{A_{\text{clean}}}.$ (1)
A robustness score of model $f$ for perturbation $\phi$, namely $R^{\phi}_{f}$, close to one indicates that the model is highly resistant to perturbation, demonstrating its ability to maintain performance in challenging conditions.
@@ IV Experimental Setup
This section presents the experimental setup used to evaluate the performance of the models. We describe the hardware and software configurations, along with the specific parameters used for training and inference. Additionally, we outline the data augmentation techniques and noise simulations applied to mimic real-world conditions in satellite-based EO tasks. The experiments focus on comparing traditional CNNs, ResNet-based models, and pre-trained ViT models, including MobileViTV2 and EfficientViT-M2, to assess their effectiveness in achieving high accuracy, computational efficiency, and robustness under noisy conditions.
IV-A Dataset and ML Model Configuration
First, to ensure a balanced and comprehensive evaluation of the model’s performance, the dataset was split into training and testing sets with a 70/30 ratio. For model optimization, different strategies have been employed based on the model architecture. The CNN and ResNet models were optimized using the Stochastic Gradient Descent (SGD) optimizer [45], configured with a learning rate of 1e-3, momentum of 0.9, and a weight decay of 5e-4. In contrast, the ViT-based models, including the pre-trained variants, have been optimized using the AdamW optimizer [46]. This setup also utilized a learning rate 1e-3 and a weight decay 5e-4, ensuring consistency across different model types. A step learning rate scheduler was applied to further enhance training efficiency with a step size of 7 epochs and a decay factor ($\gamma$) of 0.1.
The CCT [34] and SmallViT models [35] have been trained using code adapted from the ViT-PyTorch repository, which provides reliable and optimized implementations for ViT. Other pre-trained models have been sourced from Huggingface’s model hub or the official repositories in the respective papers. This approach ensured that the models were implemented consistently, with standardized architectures and preprocessing steps, allowing for a fair and direct comparison of their performance across the different experiments.222 Available codes and pre-trained models:
@@ IV-A Dataset and ML Model Configuration
IV-B Hardware setup
All experiments have been conducted on a workstation equipped with an Intel(R) Xeon(R) W-11855M CPU, operating at 3.20GHz with 12 cores and supported by 64 GB of RAM. The models have been trained and tested using an Nvidia RTX A5000 GPU with 48 GB of GDDR6 memory, providing sufficient computational resources for most of the tasks. This setup offered a robust and efficient environment for handling most models’ training and evaluation processes.
However, due to their increased complexity and size, more extensive computational resources are necessary for the two larger pre-trained ViT models (ViT-Large and ViT-Huge). These models were parallelly trained on a high-performance computing workstation with the following specifications: an Intel Xeon w9-3475X CPU featuring 36 cores and 72 threads, a maximum turbo frequency of 4.8 GHz, and an 82.5 MB cache. The system is also equipped with 512 GB of DDR5 RAM, clocked at 4800 MHz, and dual NVIDIA RTX™ 6000 Ada Generation GPUs, each with 48 GB of GDDR6 memory. This advanced hardware configuration ensured efficient training of the larger models, leveraging parallel processing and extensive memory resources.
IV-C Noisy Data for Inference
After selecting the best model with lower computational complexity and high performance, we assess its robustness during onboard satellite inference. This evaluation focuses on the model’s ability to maintain performance under noisy conditions commonly encountered in real-world satellite operations. Although the model has been trained on a clean dataset, the test data used during inference may be corrupted by noise introduced by natural events or instrument movement. This robustness analysis is critical to ensure reliable model performance in dynamic and unpredictable environments. Specifically, we introduce two types of noise into the inference test data: Gaussian noise and instrument noise represented by motion blur. For every kind of noise, we define five levels of severity, as illustrated in Fig. 1. This evaluation helps us understand how well the models can perform under varying noise levels, which is critical for reliable satellite-based IC. The noise and severity level was implemented based on the code [^3] https://github.com/hendrycks/robustness from [43].
To further assess the robustness and generalization capabilities of the selected models, we applied a series of data augmentation techniques, as illustrated in Fig. 2. The augmentations include Standard Transformed, Rand Augment [47], Random Erasing [48], Rand (Augment+Erasing), and Strong Augmentation, which combine all four augmentations designed to simulate challenging variations in the input data. These transformations introduce different levels of distortion and occlusion, such as color shifts, random erasures, and severe augmentations. By evaluating the models on these augmented datasets, we aim to determine their ability to maintain performance when faced with altered or corrupted inputs, thereby ensuring their reliability in diverse and unpredictable conditions.
Figure 1: Different noise levels with Gaussian and motion blur.
@@ V Results and Discussions
TABLE II: Summary of model computation complexity, estimated total size, training, and inference times
Models Input Size Total Parameters Size (MB) FLOPs Training (s) Inference (s)
@@
Table II provides a detailed comparison of the computational complexity of various models, with a particular emphasis on total parameters, estimated total size, floating point operations per second (FLOPs), training time, and inference time. Notably, the ViT-large and ViT-huge models stand out for their substantial complexity. They are characterized by high parameter counts, large memory footprints, and extensive computational operations (FLOPs), resulting in significantly longer training and inference times. In contrast, smaller models like EfficientViT-M2 and MobileViT2, despite being part of the ViT family, demonstrate much lower complexity across all metrics. These models are designed to be more lightweight, with fewer parameters and reduced FLOPs, making them more efficient in terms of both computational and memory resources. Compared to the pre-trained ViT models, EfficientViT-M2 and MobileViT2 balance performance and efficiency, providing a viable alternative for limited computational resources or faster processing times. This comparison underscores the diversity within Transformer-based models and highlights the importance of selecting models that align with the specific needs of a given application.
@@
TABLE III: Experimental results for the best model’s performance
@@
Based on the best run for each model in Table III, the summarized experimental results identify MobileViTV2, EfficientViT-M2, and SwinTransformer as the top-performing models. MobileViTV2 (Apple) achieves the highest overall performance, with a perfect training accuracy of 100%, leading to test accuracy, precision, and recall, each at 99.09%. SwinTransformer follows closely, boasting a training accuracy of 99.8% and a test accuracy, precision, and recall of 98.83%. EfficientViT-M2 ranks third, delivering impressive results with a training accuracy of 99.99% and test accuracy, precision, and recall, all at 98.76%. When comparing the top four large pre-trained models, the SwinTransformer notably outperforms its counterparts despite having a smaller parameter count of 86 million. It surpasses the ViT-base, ViT-large, and ViT-huge models, which do not achieve better results despite their larger parameter sizes. This highlights the effectiveness of the Swin Transformer’s design, which utilizes a hierarchical ViT architecture. Unlike conventional ViT, which relies on a global attention mechanism across the entire image, the SwinTransformer divides images into non-overlapping patches and hierarchically processes them. This method allows it to handle images at various scales and resolutions efficiently, resulting in better generalization and improved task performance. The superior performance of the SwinTransformer, even with fewer parameters, underscores the advantages of its architecture in effectively capturing both local and global image features, making it a more efficient and effective model compared to the larger, conventional attention-based ViT models.
Figure 4: Statistical comparison for power consumption during inference.
Fig. 4 compares power consumption in W during the inference on GPU. Among the pretrained ViT-based models, EfficientViT-M2 stands out for its significantly lower power consumption, registering at 29.04 $\pm$ 0.96 W. This makes it the most energy-efficient choice compared to other ViT-based models, such as MobileViTV2, which consumes 79.23 $\pm$ 1.45 W, and even more power-intensive models like ViT-huge, which uses 552.96 $\pm$ 6.19 W. The EfficientViT-M2 model also compares favorably with models trained from scratch, such as CNNs and ResNets, which exhibit power consumption of 15.42 $\pm$ 0.52 W and 24.63 $\pm$ 1.63 W, respectively. This demonstrates that EfficientViT-M2 achieves a comparable level of power efficiency during inference despite its more complex architecture and superior performance. This balance of computational power and efficiency makes it an optimal choice for onboard satellite processing, where energy consumption is a critical constraint.
@@
From the results presented in Tables II, and III, and Fig. 3, and 4 we confirm that MobileViT and EfficientViT are the two best models for IC. These models excel due to their lower computational complexity, shorter training and inference times, and superior performance across all evaluation metrics. Notably, compared to state-of-the-art models, as shown in Table IV, the exceptional performance of EfficientViT-M2 and MobileViTV2 becomes evident. In the EuroSAT IC task, MobileViTV2 achieves the highest scores across all evaluation metrics, outperforming all other models in the comparison. EfficientViT-M2 also performs exceptionally well, securing the second-highest results. These findings highlight the effectiveness of both EfficientViT-M2 and MobileViTV2 in achieving top-tier performance with relatively lower computational complexity and faster training and inference times, as observed in our experiment results. This confirms that these two models are currently the best choices for high-performance IC tasks, providing a significant advantage over existing methods.
@@
Additionally, Fig. 5 shows the confusion matrices for EfficientViT-M2 and MobileViTV2, highlighting their exceptional performance in IC. Both models exhibit vital accuracy across all classes, with most predictions aligning closely with the true labels. In the confusion matrix for EfficientViT-M2, we observe that classes like AnnualCrop, Forest, PermanentCrop, and SeaLake are classified with nearly perfect accuracy, showing minimal misclassification. Similarly, the confusion matrix for MobileViTV2 shows even fewer errors, with several classes like Forest, PermanentCrop, and SeaLake being classified almost flawlessly. A few minor misclassifications are observed in both models, particularly in categories like Pasture and River, but these are relatively small compared to the overall high performance. These matrices underscore the effectiveness of both models in handling complex IC tasks, with MobileViTV2 slightly outperforming EfficientViT-M2 in terms of fewer misclassifications, further affirming its position as the top-performing model.
@@
TABLE IV: Performance comparison with the-state-of-the-art models (Best run)
@@
Figure 6: MobileViT (Top) and EfficientViT (Bottom) robustness with Gaussian noisy inference data.
The experiment results, as presented in Tables V and VI, reveal that applying data augmentation techniques during training and inference generally leads to a slight reduction in model performance when evaluated on clean test data. MobileViTV2’s and EfficientViT-M2’s baseline performance—achieved without augmentation—yield the highest accuracy. However, among the various augmentation techniques tested, RandErasing for both models demonstrates the most effective results, providing performance metrics (98.73% for MobileViTV2 and 98.22% for EfficientViT-M2) that are closest to the baseline. This indicates that RandErasing introduces sufficient variability to enhance robustness without significantly compromising the models’ ability to perform well on clean data. Although there is a slight decrease compared to the baseline, using RandErasing offers a good balance, maintaining high accuracy while preparing the models to handle real-world scenarios with noisy inputs.
@@
Furthermore, the feasibility of deploying the EfficientViT-M2 model in practical applications, mainly OSS, is underscored by its relatively low computational complexity and compact size. With an estimated total size of 38.19 MB and a computational complexity of 203.53 MFLOPs, EfficientViT-M2 offers a significant advantage over earlier models used in pioneering satellite missions. For instance, the $\Phi$-Sat-1 mission [14], which utilized the CloudScout model for image segmentation, required approximately 4.67 GFLOPs and had an estimated total size of 123.98 MB. In contrast, EfficientViT-M2 requires substantially less computational power and storage, making it a more efficient choice for similar tasks. Additionally, compared to the $\Phi$-Sat-2 mission [15], which employed a convolutional autoencoder with 81.2 MFLOPs for vision processing on an Intel Movidius Myriad 2 VPU, EfficientViT-M2’s slightly higher complexity is justified by its enhanced robustness and performance capabilities. While the $\Phi$-Sat-2 model took nearly 11 seconds to compress an image with a compression ratio of 8 on a VPU, EfficientViT-M2’s processing requirements are more aligned with modern GPU capabilities, completing complex tasks much faster and with greater accuracy. These factors make EfficientViT-M2 a powerful model for IC and a practical and feasible option for real-world deployment in resource-constrained environments like satellite missions.